# Enhanced Monocular Visual Odometry with AR Poses and Integrated INS-GPS for Robust Localization in Urban Environments

### v1

@@ Abstract
This paper introduces a cost-effective localization system combining monocular visual odometry (VO), augmented reality (AR) poses, and integrated INS-GPS data. We address monocular VO’s scale factor issues using AR poses and enhance accuracy with INS and GPS data, filtered through an Extended Kalman Filter (EKF). Our approach, tested using manually annotated trajectories from Google Street View, achieves an RMSE of 1.529 meters over a 1 km track. Future work will focus on real-time mobile implementation and further integration of visual-inertial odometry for robust localization. This method offers lane-level accuracy with minimal hardware, making advanced navigation more accessible.
@@ I Introduction
In recent years, many cars have released with SAE Level 2 autonomous driving features. The price tag for a car with these new features is still out of reach of many people, including those who may need them most. One of the major contributors to the price of these vehicles is the cost of their sensors. Currently, a full suite of sensors for a level 2 autonomous vehicle can cost anywhere from \$8,000-10,000[1]. As we begin to progress towards higher levels of autonomy, the suite of sensors needed will only increase in complexity, and in turn, price. Additionally, it has been shown that traditionally GPS is not accurate enough to be used as a localization method for autonomous driving and this inaccuracy only increases exponentially in an urban environment. With our approach we aim to create a level of accuracy that is comparable to that of much more expensive sensors at a fraction of the cost.
In this paper we propose a way to exploit this ‘urban jungle’ and improve greatly upon the accuracy of a traditional GPS using sensors found on a single smartphone. Our method is based around using Google Street View’s panoramic images and comparing them to those being taken on board with our monocular smartphone camera. We chose to use Street View because it contains data from a large portion of the world and this data gives us both a high resolution image with accurate associated location metadata. Additionally, anyone around the world with an internet connection can access Street View. We also log measurements from the IMU and Apple’s ARKit Framework simultaneously with the images recorded from our smartphone.
Our approach involves extracting features from our recorded frames, as well as from the panoramas taken from Google Street View. Once we have these features, we calculate the relative pose between our frame and the surrounding panoramas. We then use the poses in tandem with the coordinates of each panorama provided in the metadata to triangulate our estimated position. Additionally, we aim to use the IMU measurements alongside our image data to compute our Visual-Inertial Odometry module and combine it with our estimated location in order to reduce noise.
@@ II Related Work
Previous work exists on visual global localization, including work specifically on the Google Street View dataset. Agarwal et al. [9] used optical-flow to create tracks of 3D points and performs a semi-global bundle adjustment on these tracks. More recently, Yu et al. [8] took a more limited approach by only performing a local bundle adjustment on the current set of images, not taking advantage of any odometry data. This approach showed centimeter-level accuracy, however it also made use of depth data that we found unreliable in many areas, and two roof-mounted sideways cameras that faces street facades. Our approach builds off of this previous work from [9] and [8] but uses only a single front-facing phone camera and combines additional odometry data directly from the phone.
@@ III Method
Figure 1: Diagram illustrating our system pipeline
In this section, we describe our pipeline for global localization. The pipeline was implemented in an offline manner, but at each timestep we only look at current inputs and the estimated vehicle pose, discarding all other past inputs. Although we do perform bundle adjustment, our approach does not involve SLAM in its typical formulation since we do not perform mapping. Instead, we seek to only compute a global localization through feature matching and local bundle adjustment in combination with Visual-Inertial Odometry. The diagram in Figure 1 illustrates the important steps of our pipeline.
@@ III-A Google Street View Data
The depth map captures data from laser scans or using optical flow methods, and stores only the dominant scene features such as buildings and roads [7]. However, as seen in Figure 2, in dense urban environments where we need complete representation of static scene elements, the depth map is unreliable in generating that data.
@@
To perform feature extraction and matching on the panoramas, we must generate rectilinear images from the equirectangular panoramas. Google provides an API to pull rectilinear images at specified panorama locations given the heading and pitch angles. Instead, we store a database of panoramas we intend to query as equirectangular projections of the full 360° field-of-view, and extract the rectilinear regions of the panorama with the precise vehicle heading, making our method more robust and fast. We create a “virtual camera” and perform a back-projection technique to project the panorama into a standard rectilinear view. Provided a heading and field of view, we can compute the corresponding pixel coordinates on the rectilinear image from the panorama. With the center elevation and azimuthal angles of the image being $(\theta_{0},\phi_{0})$ and the angular distance from rectilinear pixel coordinates $(x,y)$ to panorama coordinates $(\theta,\phi)$ defined as $c$, the transformation equations described in (1) are used to generate rectilinear images with a defined focal length and FoV.
@@
This virtual camera has the same camera intrinsics as our vehicle camera, maintaining the same focal length, horizontal and vertical field-of-view, and resolution, although we assume the principal point of the camera is the image center. Consistency in the camera intrinsics will simplify future computations for 3D point estimation and pose estimation knowing that the focal lengths are the same.
@@ III-B Feature Matching
For each frame in our video stream, we find nearby Street View panoramas using a KD-tree. For our image frame and one of these nearby panoramas, we use SIFT to extract key points in both images, shown in 3 [2]. With these SIFT key points, we then use FLANN based matching to find matching key points between our image and the panorama, similarly shown in 3 [3]. From this example set of matches, we notice no clear set of matching features between the two images. Clearly, a large portion of the computed matches in this example are incorrect and as a result should be discarded. These incorrect matches, referred to onwards as outliers, negatively affects our ability to find the essential matrix that describes the epipolar geometry between our captured image and the nearby panorama, which leads to inaccurate localization.
@@
To effectively address the outlier issue, we apply a filtering technique to our FLANN matches using second order descriptors known as Virtual Line Descriptors (VLD) [4]. To define a VLD, for each pair of key points in an image we compute a number of separate SIFT-like descriptors for the region of the image between the two points, shown in 4. Together, this collection of descriptors act as our second order descriptor, or VLD. With this definition, we can then also compute the difference between two VLDs as the pairwise difference between corresponding disks in their gradient histograms as well as main orientations [4].
@@
Now, we consider a pair of matches in our image and the panorama, giving us two points, or one VLD, in each of the two images. Then, we find the difference between these two VLDs, and if this difference is below a maximum threshold we consider this pair of matches to be consistent [4].
The K-VLD filtering algorithm uses this idea of consistent matches, and for any primary match between our image and the panorama, we look one by one at all the nearby matches, and for each of these, check if the pair of matches is consistent [4]. If there are at least K consistent matches for our primary match, it satisfies the K-VLD consistent constraint and is kept as a reliable match [4]. This process is repeated for all matches between the two images until we have only K-VLD consistent matches remaining [4].
We return to our previously computed FLANN matches in 3 with all of the outliers which we need to remove. Using the K-VLD algorithm, we can find all of the consistent VLDs, shown in 5 where these VLDs are highlighted. We then keep all matches from these consistent VLDs, leaving a much smaller set of matches, but now with almost all outliers removed, shown in 5.
@@
The main benefit of K-VLD filtering for our application is its property of removing ambiguous matches [4]. These types of matches are common in our dataset, which features large numbers of ambiguous features such as building windows. Additionally, this method removes false matches near epipolar lines, which again is well suited for our dataset with the majority of detected features being located along or near epipolar lines [4]. We perform SIFT feature extraction, FLANN matching, and K-VLD filtering using a fork of openMVG [25].
@@ III-C Visual Pose Estimation
In order to estimate the pose of the vehicle, we tested two similar but distinct strategies, both based on a local bundle adjustment. The first method relies upon known fixed panorama locations and uses a limited local bundle adjustment to estimate 3D points common between all panorama images. Then, using features common with the phone camera, we use PnP to estimate the phone camera relative to a reference panorama. The second method involves a standard local bundle adjustment procedure where we allow the known panorama locations as well as the vehicle pose to be adjusted subject to some constraints and use the final vehicle pose from this computation. We chose these two approaches because we sought to analyze whether allowing for more flexibility during bundle adjustment w.r.t camera poses would produce more accurate localization.
@@ III-D Fixed-pose Local Bundle Adjustment
With a set of panoramas $P$ and their corresponding features $F$ and matches $M$ with the vehicle frame $V$, we will take the subset of features
@@
that are present in all panoramas $P$ and vehicle frame $V$. This will give us a set $M^{*}$ containing pairs of features, where for each feature in $F_{V}$ we will have a corresponding feature $F_{P}$ in each of the panoramas, as seen in Figure 116. We can then perform a nonlinear optimization using SciPy to estimate the 3D points for each feature $F_{V}$ relative to the selected panorama by minimizing the following loss function. Unlike a typical bundle adjustment where we allow the camera poses to be adjusted, we fix these poses and only adjust the 3D point estimations while minimizing the error defined in (2).
@@
With the image coordinates and 3D global coordinates of $F_{V}$, we can perform a Perspective-n-Point pose estimation to best estimate the pose of the camera relative to the 3D points to best minimize the reprojection error. The overall problem of 3D point estimation can be seen in Figure 7, where we estimate the positions of the 3D object points when each of the corresponding feature points in each camera view is known.
@@ III-E Standard Local Bundle Adjustment
As before, we have a set of panoramas $P$ and their corresponding features $F$ and matches $M$ with the vehicle frame $V$. Instead of taking the intersection of features
@@
that are present in all panoramas $P$, we instead require only that a feature is present in at least $n$ panoramas. This will give us a set $M^{*}$ containing pairs of features, where for each feature in $F_{V}$ we will have a corresponding feature $F_{P}$ in at least $n$ panoramas.
To get the best results from the local bundle adjustment, we need to initialize the solver with an estimate of all $|M^{*}|$ 3D points. Since each of our $|M^{*}|$ 3D points is present in at least $n$ panoramas, there is not one correct initialization. Instead, we use the the 8-pt algorithm for each $(f_{v},f_{p}),f_{v}\in F_{V},f_{p}\in F_{p}$ pair to approximate the essential matrix. We then recover the relative pose transformation between the cameras and perform 3D reconstruction by triangulation. We then retrieve only the 3D points that are common to at least $n$ images.
Thus we have at least $n$ estimates of any given 3D point through our approximate reconstruction method. However, these points are in the reference frame of the respective panoramas so we must transform these coordinates into a global frame which we define as the reference frame of the panorama with highest number of feature matches. Given at least $n$ estimates of any given 3D point we pick the point with the median L2 norm as a simple form of outlier rejection.
We chose to compare two common nonlinear solvers used to for graph optimization, g2o [5] and Ceres [6]. A graph is a natural representation of the bundle adjustment problem since it allows us to relate quantities we wish to estimate such as camera poses and 3D points, with our measurements as edges that relate the quantities. Specifically, we construct a graph such that the edges represent the projection of landmark 3D points onto our camera frames. A visual representation of how the factor graph represents an objective function is seen in figure 8 courtesy of [5].
@@
We create $n$ vertices in our factor graph $G$ corresponding to our $n$ panorama locations and, for Ceres, we fix the allowable translation and rotation of these poses to some upper and lower bound. With more data on the accuracy of these panorama poses, future work may incorporate a Gaussian prior term in the error function that represents our certainty of the panorama poses and penalizes significant adjustments. This would better model our assumption of the system but we were unable to incorporate it into the Ceres solver through the Python bindings used for testing.
Next, we add an additional vertex corresponding to our phone camera pose which we initialize from our Kalman Filter or, for the first frame, our GPS coordinates. This vertex is entirely unconstrained.
For each of our $|M^{*}|$ 3D points, we add a vertex to our factor graph $G$ with the estimated position from our 3D reconstruction and attach at least $n+1$ edges to this point. At least $n$ edges are between panorama locations and the point with an additional edge connecting the smartphone camera frame. These edges contain the observed $(x,y)$ pixel that corresponds to the 3D point as well as the uncertainty of the measurement. The objective that we seek to minimize is the reprojection error of these 3D points onto the least $n+1$ cameras in their respective locations, and we do so by adjusting the positions of the 3D points and the 6-DoF poses of the cameras.
As suggested in [8] we use the Tukey Biweight function as in (3) with $c=3$ to improve convergence of our optimization problem in cases where outliers remain past the K-VLD step. We replace the squared error we are trying to minimize with the Tukey Biweight function and use g2o and Ceres to solve the graph optimization problem.
@@ III-F Sensor Fusion
Our estimate of the vehicle’s position using Google Street View might suffer from various points of failure such as occlusion of static features, lighting and weather condition and inaccurate street view data. The dependability of autonomous systems of one just one such position estimate can result in serious consequences. We address this problem by implementing a multi-sensor fusion algorithm. The recent trend in high quality and low cost smartphones has led to multiple sensors being incorporated into a small form factor. We exploit these sensors to derive a better estimate of the vehicle’s state. We explored two such methods:

### v2

@@ Abstract
This paper introduces a cost-effective localization system combining monocular visual odometry (VO), augmented reality (AR) poses, and integrated INS-GPS data. I address monocular VO’s scale factor issues using AR poses and enhance accuracy with INS and GPS data, filtered through an Extended Kalman Filter (EKF). My approach, tested using manually annotated trajectories from Google Street View, achieves an RMSE of 1.529 meters over a 1 km track. Future work will focus on real-time mobile implementation and further integration of visual-inertial odometry for robust localization. This method offers lane-level accuracy with minimal hardware, making advanced navigation more accessible.
@@ I Introduction
In recent years, many cars have released with SAE Level 2 autonomous driving features. The price tag for a car with these new features is still out of reach of many people, including those who may need them most. One of the major contributors to the price of these vehicles is the cost of their sensors. Currently, a full suite of sensors for a level 2 autonomous vehicle can cost anywhere from \$8,000-10,000[1]. As I begin to progress towards higher levels of autonomy, the suite of sensors needed will only increase in complexity, and in turn, price. Additionally, it has been shown that traditionally GPS is not accurate enough to be used as a localization method for autonomous driving and this inaccuracy only increases exponentially in an urban environment. With my approach I aim to create a level of accuracy that is comparable to that of much more expensive sensors at a fraction of the cost.
In this paper I proposed a way to exploit this ‘urban jungle’ and improve greatly upon the accuracy of a traditional GPS using sensors found on a single smartphone. My method is based around using Google Street View’s panoramic images and comparing them to those being taken on board with my monocular smartphone camera. I chose to use Street View because it contains data from a large portion of the world and this data gives me both a high resolution image with accurate associated location metadata. Additionally, anyone around the world with an internet connection can access Street View. I also log measurements from the IMU and Apple’s ARKit Framework simultaneously with the images recorded from my smartphone.
My approach involves extracting features from my recorded frames, as well as from the panoramas taken from Google Street View. Once I have these features, I calculate the relative pose between my frame and the surrounding panoramas. I then use the poses in tandem with the coordinates of each panorama provided in the metadata to triangulate my estimated position. Additionally, I aim to use the IMU measurements alongside my image data to compute my Visual-Inertial Odometry module and combine it with my estimated location in order to reduce noise.
@@ II Related Work
Previous work exists on visual global localization, including work specifically on the Google Street View dataset. Agarwal et al. [9] used optical-flow to create tracks of 3D points and performs a semi-global bundle adjustment on these tracks. More recently, Yu et al. [8] took a more limited approach by only performing a local bundle adjustment on the current set of images, not taking advantage of any odometry data. This approach showed centimeter-level accuracy, however it also made use of depth data that I found unreliable in many areas, and two roof-mounted sideways cameras that faces street facades. My approach builds off of this previous work from [9] and [8] but uses only a single front-facing phone camera and combines additional odometry data directly from the phone.
@@ III Method
Figure 1: Diagram illustrating my system pipeline
In this section, I describe my pipeline for global localization. The pipeline was implemented in an offline manner, but at each timestep I only look at current inputs and the estimated vehicle pose, discarding all other past inputs. Although I do perform bundle adjustment, my approach does not involve SLAM in its typical formulation since I do not perform mapping. Instead, I seek to only compute a global localization through feature matching and local bundle adjustment in combination with Visual-Inertial Odometry. The diagram in Figure 1 illustrates the important steps of my pipeline.
@@ III-A Google Street View Data
The depth map captures data from laser scans or using optical flow methods, and stores only the dominant scene features such as buildings and roads [7]. However, as seen in Figure 2, in dense urban environments where I need complete representation of static scene elements, the depth map is unreliable in generating that data.
@@
To perform feature extraction and matching on the panoramas, I must generate rectilinear images from the equirectangular panoramas. Google provides an API to pull rectilinear images at specified panorama locations given the heading and pitch angles. Instead, I store a database of panoramas I intend to query as equirectangular projections of the full 360° field-of-view, and extract the rectilinear regions of the panorama with the precise vehicle heading, making my method more robust and fast. I create a “virtual camera” and perform a back-projection technique to project the panorama into a standard rectilinear view. Provided a heading and field of view, I can compute the corresponding pixel coordinates on the rectilinear image from the panorama. With the center elevation and azimuthal angles of the image being $(\theta_{0},\phi_{0})$ and the angular distance from rectilinear pixel coordinates $(x,y)$ to panorama coordinates $(\theta,\phi)$ defined as $c$, the transformation equations described in (1) are used to generate rectilinear images with a defined focal length and FoV.
@@
This virtual camera has the same camera intrinsics as my vehicle camera, maintaining the same focal length, horizontal and vertical field-of-view, and resolution, although I assume the principal point of the camera is the image center. Consistency in the camera intrinsics will simplify future computations for 3D point estimation and pose estimation knowing that the focal lengths are the same.
@@ III-B Feature Matching
For each frame in my video stream, I find nearby Street View panoramas using a KD-tree. For my image frame and one of these nearby panoramas, I use SIFT to extract key points in both images, shown in 3 [2]. With these SIFT key points, I then use FLANN based matching to find matching key points between my image and the panorama, similarly shown in 3 [3]. From this example set of matches, I notice no clear set of matching features between the two images. Clearly, a large portion of the computed matches in this example are incorrect and as a result should be discarded. These incorrect matches, referred to onwards as outliers, negatively affects my ability to find the essential matrix that describes the epipolar geometry between my captured image and the nearby panorama, which leads to inaccurate localization.
@@
To effectively address the outlier issue, I apply a filtering technique to my FLANN matches using second order descriptors known as Virtual Line Descriptors (VLD) [4]. To define a VLD, for each pair of key points in an image I compute a number of separate SIFT-like descriptors for the region of the image between the two points, shown in 4. Together, this collection of descriptors act as my second order descriptor, or VLD. With this definition, I can then also compute the difference between two VLDs as the pairwise difference between corresponding disks in their gradient histograms as well as main orientations [4].
@@
Now, I consider a pair of matches in my image and the panorama, giving me two points, or one VLD, in each of the two images. Then, I find the difference between these two VLDs, and if this difference is below a maximum threshold I consider this pair of matches to be consistent [4].
The K-VLD filtering algorithm uses this idea of consistent matches, and for any primary match between my image and the panorama, I look one by one at all the nearby matches, and for each of these, check if the pair of matches is consistent [4]. If there are at least K consistent matches for my primary match, it satisfies the K-VLD consistent constraint and is kept as a reliable match [4]. This process is repeated for all matches between the two images until I have only K-VLD consistent matches remaining [4].
I return to my previously computed FLANN matches in 3 with all of the outliers which I need to remove. Using the K-VLD algorithm, I can find all of the consistent VLDs, shown in 5 where these VLDs are highlighted. I then keep all matches from these consistent VLDs, leaving a much smaller set of matches, but now with almost all outliers removed, shown in 5.
@@
The main benefit of K-VLD filtering for my application is its property of removing ambiguous matches [4]. These types of matches are common in my dataset, which features large numbers of ambiguous features such as building windows. Additionally, this method removes false matches near epipolar lines, which again is well suited for my dataset with the majority of detected features being located along or near epipolar lines [4]. I perform SIFT feature extraction, FLANN matching, and K-VLD filtering using a fork of openMVG [25].
@@ III-C Visual Pose Estimation
In order to estimate the pose of the vehicle, I tested two similar but distinct strategies, both based on a local bundle adjustment. The first method relies upon known fixed panorama locations and uses a limited local bundle adjustment to estimate 3D points common between all panorama images. Then, using features common with the phone camera, I use PnP to estimate the phone camera relative to a reference panorama. The second method involves a standard local bundle adjustment procedure where I allow the known panorama locations as well as the vehicle pose to be adjusted subject to some constraints and use the final vehicle pose from this computation. I chose these two approaches because I sought to analyze whether allowing for more flexibility during bundle adjustment w.r.t camera poses would produce more accurate localization.
@@ III-D Fixed-pose Local Bundle Adjustment
With a set of panoramas $P$ and their corresponding features $F$ and matches $M$ with the vehicle frame $V$, I will take the subset of features
@@
that are present in all panoramas $P$ and vehicle frame $V$. This will give me a set $M^{*}$ containing pairs of features, where for each feature in $F_{V}$ I will have a corresponding feature $F_{P}$ in each of the panoramas, as seen in Figure 116. I can then perform a nonlinear optimization using SciPy to estimate the 3D points for each feature $F_{V}$ relative to the selected panorama by minimizing the following loss function. Unlike a typical bundle adjustment where I allow the camera poses to be adjusted, I fix these poses and only adjust the 3D point estimations while minimizing the error defined in (2).
@@
With the image coordinates and 3D global coordinates of $F_{V}$, I can perform a Perspective-n-Point pose estimation to best estimate the pose of the camera relative to the 3D points to best minimize the reprojection error. The overall problem of 3D point estimation can be seen in Figure 7, where I estimate the positions of the 3D object points when each of the corresponding feature points in each camera view is known.
@@ III-E Standard Local Bundle Adjustment
As before, I have a set of panoramas $P$ and their corresponding features $F$ and matches $M$ with the vehicle frame $V$. Instead of taking the intersection of features
@@
that are present in all panoramas $P$, I instead require only that a feature is present in at least $n$ panoramas. This will give me a set $M^{*}$ containing pairs of features, where for each feature in $F_{V}$ I will have a corresponding feature $F_{P}$ in at least $n$ panoramas.
To get the best results from the local bundle adjustment, I need to initialize the solver with an estimate of all $|M^{*}|$ 3D points. Since each of my $|M^{*}|$ 3D points is present in at least $n$ panoramas, there is not one correct initialization. Instead, I use the the 8-pt algorithm for each $(f_{v},f_{p}),f_{v}\in F_{V},f_{p}\in F_{p}$ pair to approximate the essential matrix. I then recover the relative pose transformation between the cameras and perform 3D reconstruction by triangulation. I then retrieve only the 3D points that are common to at least $n$ images.
Thus I have at least $n$ estimates of any given 3D point through my approximate reconstruction method. However, these points are in the reference frame of the respective panoramas so I must transform these coordinates into a global frame which I define as the reference frame of the panorama with highest number of feature matches. Given at least $n$ estimates of any given 3D point I pick the point with the median L2 norm as a simple form of outlier rejection.
I chose to compare two common nonlinear solvers used to for graph optimization, g2o [5] and Ceres [6]. A graph is a natural representation of the bundle adjustment problem since it allows me to relate quantities I wish to estimate such as camera poses and 3D points, with my measurements as edges that relate the quantities. Specifically, I construct a graph such that the edges represent the projection of landmark 3D points onto my camera frames. A visual representation of how the factor graph represents an objective function is seen in figure 8 courtesy of [5].
@@
I create $n$ vertices in my factor graph $G$ corresponding to my $n$ panorama locations and, for Ceres, I fix the allowable translation and rotation of these poses to some upper and lower bound. With more data on the accuracy of these panorama poses, future work may incorporate a Gaussian prior term in the error function that represents my certainty of the panorama poses and penalizes significant adjustments. This would better model my assumption of the system but I were unable to incorporate it into the Ceres solver through the Python bindings used for testing.
Next, I add an additional vertex corresponding to my phone camera pose which I initialize from my Kalman Filter or, for the first frame, my GPS coordinates. This vertex is entirely unconstrained.
For each of my $|M^{*}|$ 3D points, I add a vertex to my factor graph $G$ with the estimated position from my 3D reconstruction and attach at least $n+1$ edges to this point. At least $n$ edges are between panorama locations and the point with an additional edge connecting the smartphone camera frame. These edges contain the observed $(x,y)$ pixel that corresponds to the 3D point as well as the uncertainty of the measurement. The objective that I seek to minimize is the reprojection error of these 3D points onto the least $n+1$ cameras in their respective locations, and I do so by adjusting the positions of the 3D points and the 6-DoF poses of the cameras.
As suggested in [8] I use the Tukey Biweight function as in (3) with $c=3$ to improve convergence of my optimization problem in cases where outliers remain past the K-VLD step. I replace the squared error I are trying to minimize with the Tukey Biweight function and use g2o and Ceres to solve the graph optimization problem.
@@ III-F Sensor Fusion
my estimate of the vehicle’s position using Google Street View might suffer from various points of failure such as occlusion of static features, lighting and weather condition and inaccurate street view data. The dependability of autonomous systems of one just one such position estimate can result in serious consequences. I address this problem by implementing a multi-sensor fusion algorithm. The recent trend in high quality and low cost smartphones has led to multiple sensors being incorporated into a small form factor. I exploit these sensors to derive a better estimate of the vehicle’s state. I explored two such methods: